OpenStudy (anonymous):

y=ln e^x/e^x-1 Find the first derivative...show your work

OpenStudy (anonymous):

not clear

OpenStudy (anonymous):

is it \[\log(\frac{e^x}{e^x-1})\] or \[\frac{\log(e^x)}{e^x-1}\]

OpenStudy (anonymous):

\[\ln \ \frac{e^x }{ e^x-1 }\]

OpenStudy (anonymous):

find the derivative of that